# Dissipative particle dynamics for coarse-grained models

### v1

@@ Abstract
We develop a computational method based on Dissipative Particle Dynamics (DPD) that introduces solvent hydrodynamic interactions to coarse-grained models of solutes, such as ions, molecules, or polymers. DPD-solvent (DPDS) is a fully off-lattice method that allows straightforward incorporation of hydrodynamics at desired solvent viscosity, compressibility and solute diffusivity with any particle-based solute model. Solutes interact with the solvent only through the DPD thermostat, which ensures that the equilibrium properties of the solute system are not affected by the introduction of the DPD solvent. Thus, DPDS can be used as a replacement for traditional molecular dynamics thermostats such as Nosé-Hoover and Langevin. We demonstrate the applicability of DPDS on the case of polymer dynamics and electroosmotic flow through a nanopore. The method should be broadly useful as means to introduce hydrodynamics to existing coarse-grained models of molecules and soft materials.
@@
To achieve this, the coarse-grained system can be coupled to a mesh, typically a cubic grid, and hydrodynamics is solved using this mesh. A few examples of such approaches include Multi-Particle Collision dynamics (MPC) or Stochastic Rotation Dynamics Malevanets and Kapral (1999); Gompper et al. (2009), Lattice Boltzmann (LB) methods Krueger et al. (2016), and Fluid Particle Dynamics (FPD) Tanaka and Araki (2000). However, coupling of the continuous system to a mesh can introduce artifacts and increase propagation errors. MPC does not conserve angular momentum (or breaks time-reversal symmetry) and suffers from artificially high compressibility Gompper et al. (2009). LB allows embedding of large colloids, but embedding of particles smaller than the lattice size, e.g., polymers, suffers from lattice discretization errors and mixing molecular dynamics with LB propagation leads to energy conservation issues Dünweg and Ladd (2009). FPD directly solves the Navier–Stokes equations but can only embed objects at least a few lattice sites in size, which makes the method suitable for nanoparticles and colloids Furukawa et al. (2018), but is computationally inefficient for small molecules, ions or polymers.
@@
Among off-lattice approaches the most straightforward way to partially introduce hydrodynamics is to apply a Galilean-invariant thermostat directly to the solute particles Soddemann et al. (2003), however, this approach requires a dense suspension and is not applicable to dilute solutions. A general off-lattice method is Dissipative Particle Dynamics (DPD) that simulates hydrodynamic interactions by representing the fluid as soft particles Hoogerbrugge and Koelman (1992); Groot and Warren (1997). DPD reproduces a desired solvent viscosity and compressibility while describing thermal fluctuations and allowing the explicit representation of different chemical components Espa ol and Warren (2017). However, simply adding DPD particles to a coarse-grained system introduces depletion effects and thus changes the equilibrium properties of the system. While it is possible to model a variety of multicomponent systems by modifying the conservative interactions between DPD particles Pagonabarraga and Frenkel (2001); Groot and Rabone (2001), this approach requires careful parametrization; altering conservative interactions does not merely change the equilibrium properties but also transport behavior such as local viscosity, compressibility and diffusivity. Furthermore, point charges cannot be included due to the DPD soft repulsive potentials Groot (2003).
@@
Here we propose an alternative off-lattice method, named DPD solvent (DPDS), that incorporates hydrodynamic interactions in coarse-grained models of solutes. The premise of the method is overlay the DPD fluid at desired viscosity and compressibility with the solute system. The DPD solvent interacts with the solute system only via the DPD thermostat. Thus, the equilibrium properties of the system and the DPD-solvent are fully decoupled and can be controlled separately, while the transport properties are determined by the solute–solvent thermostat coupling strength. The main advantage of this procedure over existing methods is that it avoids any lattice mapping issues. The same or a similar velocity-Verlet integrator is used to evolve configurations both for the solute system and the DPD solvent, which minimizes propagation errors due to solvent–solute coupling and significantly simplifies the implementation compared to hybrid lattice approaches.
@@
In the following, we review the DPD technique and then show how to combine it with coarse-grained systems of solutes using the DPDS approach. We demonstrate the applicability of DPDS on two typical coarse-grained models, a bead-spring polymer and a primitive model electrolyte, and show that it accurately describes hydrodynamic interactions by investigating polymer collapse, Zimm dynamics and electrokinetic flow in patterned nanochannels.
@@
where the first term describes the friction force parallel to the inter-particle vector $\mathbf{r}_{ij}=\mathbf{r}_{j}-\mathbf{r}_{i}$ with $\hat{\mathbf{r}}_{ij}=\mathbf{r}_{ij}/|\mathbf{r}_{ij}|$ the corresponding unit vector. The friction is proportional to the coefficient $\gamma$ and distance-dependent weighting function $w(r_{ij})$. The second term describes the corresponding perpendicular contribution Junghans et al. (2008), i.e., the shear friction, with coefficient $\gamma_{\perp}$ and weighting function $w_{\perp}(r_{ij})$. $\mathbf{I}$ is the identity matrix and $\hat{\mathbf{r}}_{ij}^{T}$ denotes the transpose of $\hat{\mathbf{r}}_{ij}$.
@@
where the prefactors are determined by the desired temperature $T$ of the system, $dW_{ij}$ is the independent increment of a Wiener process and $d{\mathbf{W}}_{ij}$ the vector of independent Wiener processes. For a finite time-step, $\frac{\Delta W_{ij}}{\Delta t}=\zeta_{ij}\Delta t^{-1/2}$, and $\frac{\Delta{\mathbf{W}}_{ij}}{\Delta t}=\bm{\zeta}_{ij}\Delta t^{-1/2}$ where $\zeta_{ij}$ is a symmetric random variable with zero mean and unit variance and $\bm{\zeta}_{ij}$ is the corresponding random vector.
@@
with $m$ the mass of the DPD particles. These equations are usually evolved using the velocity-Verlet integrator Groot and Warren (1997); Pagonabarraga et al. (1998) and the DPD model is implemented in various open-source molecular dynamics (MD) packages such as LAMMPS, EPResSo and HOOMD-blue. Since the dissipative and random forces act only on relative velocities, the integrator preserves Galilean invariance and conserves momentum, implying that the behavior of the fluid follows Navier-Stokes hydrodynamics on sufficiently large length scales.
When applied to molecular simulations, the main advantage of DPD compared to Langevin dynamics or finite element hydrodynamic solvers is that the DPD method simultaneously samples the correct canonical $(N,V,T)$ Gibbs-Boltzmann ensemble, preserves hydrodynamics and can capture chemical specificity. Consequently, the method has been widely used to simulate hydrodynamic interactions in various systems including colloidal suspensions, blood, phase-separating fluids, polymer solutions, electrolytes, and biological membranes; see Refs. Espa ol and Warren (2017); Santo and Neimark (2021) for extensive reviews of DPD fundamentals and applications.
In the following we use standard parameters corresponding to an aqueous solution with each DPD particle representing $N_{\textnormal{w}}=3$ water molecules: the particle density $\rho_{\textnormal{s}}=3/r_{\textnormal{c}}^{3}$, with the cutoff $r_{c}=0.646\,$nm, which is taken as the length unit $\lambda=r_{\textnormal{c}}$, and the interaction prefactor $a_{ij}=78k_{\textnormal{B}}T$ reproduces the compressibility of water at room temperature Groot and Warren (1997). To reduce the number of parameters, the dissipative weighing function $w(r_{ij})$ has the same form as the conservative repulsive force, a common choice for DPD simulations Espa ol and Warren (2017),
@@
The friction coefficient $\gamma$ determines the crossover timescale between the ballistic and diffusive regimes $\tau_{\textnormal{BD}}=m/\gamma$. The typical choice for the friction parameter is $\gamma=4.5\sqrt{k_{\textnormal{B}}Tm}/r_{\textnormal{c}}$ Groot and Warren (1997), which ensures that the dynamics is diffusive on all relevant length scales ($r\geq r_{\textnormal{c}}$) and time scales ($t\geq\tau$), where the molecular dynamics time unit $\tau$, which measures the characteristic time required to move ballistically by a distance $r_{\textnormal{c}}$, is $\tau=r_{\textnormal{c}}\sqrt{m/(k_{\textnormal{B}}T)}$. The original DPD formulation used no shear friction ($\gamma_{\perp}=0$Groot and Warren (1997), but its inclusion has been shown to improve hydrodynamic properties Junghans et al. (2008) so we use $\gamma_{\perp}=\gamma$ and $w_{\perp}(r_{ij})=w(r_{ij})$ Lauriello et al. (2021).
@@
We note that employing these standard DPD parameters contains an inherent limitation that the Schmidt number $\textnormal{Sc}=\nu/D$, which measures the ratio of momentum diffusion to mass diffusion, with $\nu$ the kinematic viscosity and $D$ the diffusion constant, is too low. $\textnormal{Sc}\sim 7$ for the DPD fluid, which is much smaller than $\textnormal{Sc}~{}\sim 370$ expected for water at room temperature ($T=298\,$K) and standard pressure Groot and Warren (1997); Lauriello et al. (2021). This discrepancy is a consequence of using a soft repulsive potential [Eq. (1)] and relatively weak friction that together allow using significantly larger time-steps, which increases the efficiency of the method by about four orders of magnitude compared to atomistic simulations Groot and Rabone (2001). However, the low Schmidt number does not appear to affect transport properties of polymers as long as momentum diffusion is faster than mass diffusion ($\textnormal{Sc}>1$Spenley (2000). Moreover, for the majority of nanoscale systems hydrodynamic flows occur in the low Reynolds number regime ($\textnormal{Re}<1$) in which case inertia is irrelevant and thus an incorrect Schmidt number is not expected to affect any observables. If required, high Schmidt numbers can be obtained by increasing the friction coefficient $\gamma$ or modifying the weighing function $w(r_{ij})$, at a cost of efficiency Krafnick and Garc a (2015); Lauriello et al. (2021), or by using the Lowe thermostat Lowe, C. P. (1999). We stress that the proposed DPDS method is general and can be applied to any DPD parameters that model higher Schmidt numbers or non-aqueous solutions.
@@
The simulation time scale $\tau$ is determined by setting the desired dynamic viscosity $\eta$ of the fluid. The viscosity of the DPD fluid can be measured via self-diffusivity, the decay of the stress autocorrelation function or the Poiseuille profile Boromand et al. (2015); Lauriello et al. (2021). At the standard parameters used the viscosity obtained by fitting the Poiseuille profile is $\eta=(2.31\pm 0.05)k_{\textnormal{B}}T\tau r_{\textnormal{c}}^{-3}$ (see Sec. Nanochannel flow), which is consistent with the value obtained by the decay of the stress autocorrelation function ($\nu=0.748k_{\textnormal{B}}T\tau r_{\textnormal{c}}^{-3}$, at $\rho=3r_{\textnormal{c}}^{-3}$auriello et al. (2021). To model the dynamic viscosity of water, $\eta=10^{-3}\,\textnormal{Pa}\cdot\textnormal{s}$, at room temperature, $T=298$ K, the MD simulation time scale is thus $\tau=29\,$ps.  [^1] Using this timescale we can determine the mass of the particles and find $m\approx 8.310^{-24}\,$kg, which is about two orders of magnitude larger than the mass of three water molecules. This is a consequence of the low Schmidt number at the standard DPD parameters and the requirement to model the desired dynamic viscosity $\eta$.
@@ II DPD solvent
(ii) The system interacts with the solvent via the DPD thermostat whose coupling is set to reproduce the desired diffusion constant of the solutes.
Specifically, a system containing $N_{\textnormal{s}}$ solute particles is described by interactions that define the potential energy $V_{\textnormal{s}}(\mathbf{r}^{N_{\textnormal{s}}}_{\textnormal{s}})$ depending on the positions $\mathbf{r}^{N_{\textnormal{s}}}_{\textnormal{s}}$ of these solute particles. To preserve the equilibrium distribution of configurational states, the DPD fluid is coupled to the solutes only via the random and dissipative forces [Eqs. (2) and (3)] and the solute–DPD coupling is determined by the strength $\gamma_{\textnormal{s}}$ and the weighting function $w_{\textnormal{s}}(r_{ij})$. To simplify the implementation and reduce the number of parameter $w_{\textnormal{s}}(r_{ij})$ has the same form as between DPD particles, Eq. (5), but with a different cutoff value $r_{\textnormal{s}}$,
@@
The diffusion constant $D=D(r_{\textnormal{s}},\gamma_{\textnormal{s}})$ measured via the mean-squared displacement is shown in Figure 1. The diffusion constant scaling can also be estimated analytically. Strong coupling, $\gamma_{\textnormal{s}}\to\infty$, implies the solvent is rigidly coupled to the solute within range $r_{\textnormal{s}}$, thus the solute behaves as a solid sphere with radius $r_{\textnormal{s}}$ whose diffusivity is given by the Stokes-Einstein relation, $D=k_{\textnormal{B}}T/(6\pi\eta r_{\textnormal{s}})$. Conversely, in the weak coupling regime, the diffusion constant can be calculated by mapping the thermostat forces to a Langevin equation Groot and Warren (1997), which yields the diffusion constant $D=15k_{\textnormal{B}}T/2\pi\gamma_{\textnormal{s}}\rho r_{\textnormal{s}}^{3}$ (note that this expression is a factor 3 smaller than the result in Ref. Groot and Warren (1997) due to inclusion of shear forces, $\gamma_{\perp}=\gamma$). These two limits imply the following functional form for the diffusion constant of the solutes,
@@
The diffusion constant of a typical small solute such as an ion or a small molecule is $D\approx 1~{}\textnormal{nm}^{2}/\textnormal{ns}$, which is $D\approx 0.078\lambda^{2}/\tau$ in the standard DPD units ($\lambda=0.646~{}\textnormal{nm}$, $\tau=0.029~{}\textnormal{ns}$). Our simulation data confirm that chosing the mass of the solute particles the same as the mass of the DPD particles the dynamics are already fully in the diffusive regime at $\langle r^{2}\rangle=\lambda^{2}$ (Fig. 3). As expected, the dynamics do not depend on specific $\gamma_{\textnormal{s}}$ and $r_{\textnormal{s}}$ as long as $D$ is kept constant. The transition from ballistic to diffusive dynamics occurs on the lengtscale $r_{\textnormal{BD}}\sim\sqrt{D\tau}\approx 0.28\lambda$. Therefore, we conclude that the standard DPD friction $\gamma=4.5k_{\textnormal{B}}T\tau r_{\textnormal{c}}^{-2}$ models the appropriate diffusive dynamics on the relevant lengthscales ($r\geq\lambda$).

### v2

@@ Abstract
We develop a computational method based on Dissipative Particle Dynamics (DPD) that introduces solvent hydrodynamic interactions to coarse-grained models of solutes, such as ions, molecules, or polymers. DPD-solvent (DPDS) is a fully off-lattice method that allows straightforward incorporation of hydrodynamics at desired solvent viscosity, compressibility, and solute diffusivity with any particle-based solute model. Solutes interact with the solvent only through the DPD thermostat, which ensures that the equilibrium properties of the solute system are not affected by the introduction of the DPD solvent, while the thermostat coupling strength sets the desired solute diffusivity. Thus, DPDS can be used as a replacement for traditional molecular dynamics thermostats such as Nosé-Hoover and Langevin. We demonstrate the applicability of DPDS in the case of polymer dynamics and electroosmotic flow through a nanopore. The method should be broadly useful as a means to introduce hydrodynamic interactions to existing coarse-grained models of solutes and soft materials.
@@
To achieve this, the coarse-grained system can be coupled to a mesh, typically a cubic grid, and hydrodynamics is solved using this mesh. A few examples of such approaches include Multi-Particle Collision dynamics (MPC) or Stochastic Rotation Dynamics Malevanets and Kapral (1999); Gompper et al. (2009), Lattice Boltzmann (LB) methods Krueger et al. (2016), and Fluid Particle Dynamics (FPD) Tanaka and Araki (2000). However, coupling of the continuous system to a mesh can introduce artifacts and increase propagation errors. MPC does not conserve angular momentum (or breaks time-reversal symmetry) and suffers from artificially high compressibility Gompper et al. (2009). LB allows embedding of large colloids, but embedding of particles smaller than the lattice size, e.g., polymers, suffers from lattice discretization errors and mixing molecular dynamics with LB propagation leads to energy conservation issues Dünweg and Ladd (2009). FPD directly solves the Navier–Stokes equations but can only embed objects at least a few lattice sites in size, which makes the method suitable for nanoparticles and colloids Furukawa, Tateno, and Tanaka (2018), but is computationally inefficient for small molecules, ions or polymers.
@@
Among off-lattice approaches the most straightforward way to partially introduce hydrodynamics is to apply a Galilean-invariant thermostat directly to the solute particles Soddemann, Dünweg, and Kremer (2003), however, this approach requires a dense suspension and is not applicable to dilute solutions. A general off-lattice method is Dissipative Particle Dynamics (DPD) that simulates hydrodynamic interactions by representing the fluid as soft particles Hoogerbrugge and Koelman (1992); Groot and Warren (1997). DPD reproduces a desired solvent viscosity and compressibility while describing thermal fluctuations and allowing the explicit representation of different chemical components Espanol and Warren (2017). However, simply adding DPD particles to a coarse-grained system introduces solvation and depletion effects, and thus changes the equilibrium properties of the system. To overcome these problems, the solute model must be designed for a specific DPD solvent parametrization and typically the solute–solute interaction is restricted to the soft DPD repulsive interaction, which significantly limits the space of solute models. For example, point charges cannot be included due to electrostatic divergence issues when combined with the DPD soft repulsive potentials Groot (2003). While it is possible to model a variety of multicomponent systems by modifying the conservative interactions between DPD particles Pagonabarraga and Frenkel (2001); Groot and Rabone (2001); Guo, Liang, and Wang (2011), this approach requires careful parametrization; altering conservative interactions does not merely change the equilibrium properties but also transport behavior such as local viscosity, compressibility and diffusivity.
@@
Here we propose an alternative off-lattice method, named DPD solvent (DPDS), that incorporates hydrodynamic interactions in coarse-grained models of solutes. The premise of the method is to overlay the DPD fluid at desired viscosity and compressibility with the solute system. The DPD solvent interacts with the solute system only via the DPD thermostat. Thus, the equilibrium properties of the system and the DPD-solvent are fully decoupled and can be controlled separately, while the transport properties are determined by the solute–solvent thermostat coupling strength. The main advantage of this procedure over existing lattice methods is that it avoids any lattice mapping issues. The same or a similar velocity-Verlet integrator is used to evolve configurations both for the solute system and the DPD solvent, which minimizes propagation errors due to solvent–solute coupling and significantly simplifies the implementation compared to hybrid lattice approaches. The main advantage of DPDS over standard DPD is that DPDS can be directly used with any solute model and is not restricted to soft DPD repulsive potentials. This enables straightforward simulations of charged particles and introduction of hydrodynamic interactions to already developed solute models. Moreover, hydrodynamic interactions can be switched on/off without affecting any equilibrium properties of the solute system, allowing direct evaluation of hydrodynamic effects compared to, for example, Langevin dynamics.
@@
In the following, we review the DPD technique and then show how to combine it with coarse-grained systems of solutes using the DPDS approach. We demonstrate the applicability of DPDS on two typical coarse-grained models, a bead-spring polymer and an electrolyte solution, and show that it accurately describes hydrodynamic interactions by investigating polymer collapse, Zimm dynamics and electrokinetic flow in patterned nanochannels.
@@
where the first term describes the friction force parallel to the inter-particle vector $\mathbf{r}_{ij}=\mathbf{r}_{j}-\mathbf{r}_{i}$ with $\hat{\mathbf{r}}_{ij}=\mathbf{r}_{ij}/|\mathbf{r}_{ij}|$ the corresponding unit vector. The friction is proportional to the coefficient $\gamma$ and distance-dependent weighting function $w(r_{ij})$. The second term describes the corresponding perpendicular contribution Junghans, Praprotnik, and Kremer (2008), i.e., the shear friction, with coefficient $\gamma_{\perp}$ and weighting function $w_{\perp}(r_{ij})$. $\mathbf{I}$ is the identity matrix and $\hat{\mathbf{r}}_{ij}^{T}$ denotes the transpose of $\hat{\mathbf{r}}_{ij}$.
@@
where the prefactors are determined by the desired temperature $T$ of the system, $dW_{ij}$ is the independent increment of a Wiener process and $d{\mathbf{W}}_{ij}$ the vector of independent Wiener processes. For a finite time-step $\Delta t$, $\frac{\Delta W_{ij}}{\Delta t}=\zeta_{ij}\Delta t^{-1/2}$, and $\frac{\Delta{\mathbf{W}}_{ij}}{\Delta t}=\bm{\zeta}_{ij}\Delta t^{-1/2}$ where $\zeta_{ij}$ is a symmetric random variable with zero mean and unit variance and $\bm{\zeta}_{ij}$ is the corresponding random vector.
@@
with $m$ the mass of the DPD particles. These equations are usually evolved using the velocity-Verlet integrator Groot and Warren (1997); Pagonabarraga, Hagen, and Frenkel (1998) and the DPD model is implemented in various open-source molecular dynamics (MD) packages such as LAMMPS, EPResSo and HOOMD-blue. Since the dissipative and random forces act only on relative velocities, the integrator preserves Galilean invariance and conserves momentum, implying that the behavior of the fluid follows Navier-Stokes hydrodynamics on sufficiently large length scales.
When applied to molecular simulations, the main advantage of DPD compared to Langevin dynamics or finite element hydrodynamic solvers is that the DPD method simultaneously samples the correct canonical $(N,V,T)$ Gibbs-Boltzmann ensemble, preserves hydrodynamics and can capture chemical specificity. Consequently, the method has been widely used to simulate hydrodynamic interactions in various systems including colloidal suspensions, blood, phase-separating fluids, polymer solutions, electrolytes, and biological membranes; see Refs. Espanol and Warren (2017); Santo and Neimark (2021) for extensive reviews of DPD fundamentals and applications.
In the following we use standard parameters corresponding to an aqueous solution with each DPD particle representing $N_{\textnormal{w}}=3$ water molecules: the particle density $\rho_{\textnormal{s}}=3/r_{\textnormal{c}}^{3}$, with the cutoff $r_{c}=0.646\,$nm, which is taken as the length unit $\lambda=r_{\textnormal{c}}$, and the interaction prefactor $a_{ij}=78k_{\textnormal{B}}T$ reproduces the compressibility of water at room temperature Groot and Warren (1997). To reduce the number of parameters, the dissipative weighing function $w(r_{ij})$ has the same form as the conservative repulsive force, a common choice for DPD simulations Espanol and Warren (2017),
@@
The friction coefficient $\gamma$ determines the crossover timescale between the ballistic and diffusive regimes $\tau_{\textnormal{BD}}\sim m/\gamma$ (exact calculation is provided below). The typical choice for the friction parameter is $\gamma=4.5\sqrt{k_{\textnormal{B}}Tm}/r_{\textnormal{c}}$ Groot and Warren (1997), which ensures that the dynamics is diffusive on all relevant length scales ($r\geq r_{\textnormal{c}}$) and time scales ($t\geq\tau$), where the molecular dynamics time unit $\tau$, which measures the characteristic time required to move ballistically by a distance $r_{\textnormal{c}}$, is $\tau=r_{\textnormal{c}}\sqrt{m/(k_{\textnormal{B}}T)}$. The original DPD formulation used no shear friction ($\gamma_{\perp}=0$Groot and Warren (1997), but its inclusion has been shown to improve hydrodynamic properties Junghans, Praprotnik, and Kremer (2008) so we use $\gamma_{\perp}=\gamma$ and $w_{\perp}(r_{ij})=w(r_{ij})$ Lauriello et al. (2021).
@@
We note that employing these standard DPD parameters contains an inherent limitation that the Schmidt number $\textnormal{Sc}=\nu_{\textnormal{k}}/D$, which measures the ratio of momentum diffusion to mass diffusion, with $\nu_{\textnormal{k}}$ the kinematic viscosity and $D$ the diffusion constant, is too low. $\textnormal{Sc}\sim 7$ for the DPD fluid, which is much smaller than $\textnormal{Sc}~{}\sim 370$ expected for water at room temperature ($T=298\,$K) and standard pressure Groot and Warren (1997); Lauriello et al. (2021). This discrepancy is a consequence of using a soft repulsive potential [Eq. (1)] and relatively weak friction that together allow using significantly larger time-steps, which increases the efficiency of the method by about four orders of magnitude compared to atomistic simulations Groot and Rabone (2001). However, the low Schmidt number does not appear to affect transport properties of polymers as long as momentum diffusion is faster than mass diffusion ($\textnormal{Sc}>1$Spenley (2000). Moreover, for the majority of nanoscale systems hydrodynamic flows occur in the low Reynolds number regime ($\textnormal{Re}<1$) in which case inertia is irrelevant and thus an incorrect Schmidt number is not expected to affect any observables. If required, high Schmidt numbers can be obtained by increasing the friction coefficient $\gamma$ or modifying the weighing function $w(r_{ij})$, at a cost of efficiency Krafnick and Garcia (2015); Lauriello et al. (2021), or by using the Lowe thermostat Lowe, C. P. (1999). We stress that the proposed DPDS method is general and can be applied to any DPD parameters that model higher Schmidt numbers or non-aqueous solutions.
@@
The simulation time scale $\tau$ is determined by setting the desired dynamic viscosity $\eta$ of the fluid. The viscosity of the DPD fluid can be measured via self-diffusivity, the decay of the stress autocorrelation function or the Poiseuille profile Boromand, Jamali, and Maia (2015); Lauriello et al. (2021). At the standard parameters used the viscosity obtained by fitting the Poiseuille profile is $\eta=(2.31\pm 0.05)k_{\textnormal{B}}T\tau r_{\textnormal{c}}^{-3}$ (see Sec. Nanochannel flow), which is consistent with the value obtained by the decay of the stress autocorrelation function ($\nu_{\textnormal{k}}=0.748k_{\textnormal{B}}T\tau m^{-1}$, at $\rho=3r_{\textnormal{c}}^{-3}$auriello et al. (2021). To model the dynamic viscosity of water, $\eta=10^{-3}\,\textnormal{Pa}\cdot\textnormal{s}$, at room temperature, $T=298$ K, the MD simulation time scale is thus $\tau=29\,$ps.   [^1] Using this timescale we can determine the mass of the particles and find $m\approx 8.310^{-24}\,$kg, which is about two orders of magnitude larger than the mass of three water molecules. This is a consequence of the low Schmidt number at the standard DPD parameters and the requirement to model the desired dynamic viscosity $\eta$.
@@ II DPD solvent
(ii) The system interacts with the solvent via the DPD thermostat whose coupling is set to reproduce the desired diffusion constant of the solutes. Specifically, a system containing $N_{\textnormal{s}}$ solute particles is described by interactions that define the potential energy $V_{\textnormal{s}}(\mathbf{r}^{N_{\textnormal{s}}}_{\textnormal{s}})$ depending on the positions $\mathbf{r}^{N_{\textnormal{s}}}_{\textnormal{s}}$ of these solute particles. To preserve the equilibrium distribution of configurational states, the DPD fluid is coupled to the solutes only via the random and dissipative forces [Eqs. (2) and (3)] and the solute–DPD coupling is determined by the strength $\gamma_{\textnormal{s}}$ and the weighting function $w_{\textnormal{s}}(r_{ij})$. To simplify the implementation and reduce the number of parameters, $w_{\textnormal{s}}(r_{ij})$ has the same form as between DPD particles, Eq. (5), but with a different cutoff value $r_{\textnormal{s}}$,
@@
The diffusion constant $D=D(r_{\textnormal{s}},\gamma_{\textnormal{s}})$ measured via the mean-squared displacement is shown in Figure 1. The diffusion constant scaling can also be estimated analytically. Strong coupling, $\gamma_{\textnormal{s}}\to\infty$, implies the solvent is rigidly coupled to the solute within range $r_{\textnormal{s}}$, thus the solute behaves as a solid sphere with radius $r_{\textnormal{s}}$ whose diffusivity is given by the Stokes-Einstein relation, $D=k_{\textnormal{B}}T/(6\pi\eta r_{\textnormal{s}})$. Conversely, in the weak coupling regime, the diffusion constant can be calculated by mapping the thermostat forces to a Langevin equation Groot and Warren (1997), which yields the diffusion constant $D=15k_{\textnormal{B}}T/2\pi\gamma_{\textnormal{s}}\rho r_{\textnormal{s}}^{3}$ (note that this expression is a factor 3 smaller than the result in Ref. Groot and Warren (1997) due to inclusion of shear forces, $\gamma_{\perp,\textnormal{s}}=\gamma_{\textnormal{s}}$). These two limits imply the following functional form for the diffusion constant of the solutes,
@@
The diffusion constant of a typical small solute such as an ion or a small molecule is $D\approx 1~{}\textnormal{nm}^{2}/\textnormal{ns}$, which is $D\approx 0.078\lambda^{2}/\tau$ in the standard DPD units ($\lambda=0.646~{}\textnormal{nm}$, $\tau=0.029~{}\textnormal{ns}$). Our simulation data confirm that chosing the mass of the solute particles the same as the mass of the DPD particles the dynamics are already fully in the diffusive regime at $\langle r^{2}\rangle=\lambda^{2}$ (Fig. 3). As expected, the dynamics do not depend on specific $\gamma_{\textnormal{s}}$ and $r_{\textnormal{s}}$ as long as $D$ is kept constant. The transition from ballistic to diffusive dynamics occurs on the lengthscale $r_{\textnormal{BD}}=2\sqrt{3}D\tau/\lambda\approx 0.27\lambda$ (Fig. 3). Therefore, we conclude that the standard DPD friction $\gamma=4.5k_{\textnormal{B}}T\tau r_{\textnormal{c}}^{-2}$ models the appropriate diffusive dynamics on the relevant lengthscales ($r\geq\lambda$).